What is x(squared)- 8x - 33
Answer by nerdybill(7384)

x^2- 8x - 33
factored:
(x-11)(x+3)
.
To solve, normally it was set to zero:
x^2- 8x - 33 = 0
(x-11)(x+3) = 0
x = {-3, 11}